# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 8. Examples and remarks
depending on how $k=k_{1}+\dots+k_{t}$ compares to $p$. Such a product is furthermore robust in the sense that it is commutative in the factors provided $k\leq p$, which follows from Theorem 6.5 (vii), cf. Remark 6.6 (ii). Similarly, it follows by Proposition 7.3 (ii) that replacing any $s_{k_{j}}(E)$ by $s_{k_{j}}(E,Z)$ does not change the resulting current provided $k\leq p$. In particular, $c_{k}(E)=c_{k}(E,Z)$ for $k\leq p$. For $k>p$, we will see below that such an equality does typically not hold.
@@ 8.1. Difference between $c_{k}(E)$ and $c_{k}(E,Z)$
Example 8.1.
In general $c(E,Z)$ and $c(E)$ are different. The simplest example is when $Z=X$ and $Z^{\prime}\neq Z$. Then in general $s(E)$ may have a residue term, whereas $s(E,Z)=\omega$ and hence $c(E,Z)=c(E_{0})$. ∎
@@
For instance, if $\tau\colon{\mathbb{C}}\to{\mathbb{C}}^{2}$, $t\mapsto(t,0)$, then ${\bf s}_{1}(\tau^{*}E)=[t=0]$. Hence, by Theorem 6.5 (iii), ${\bf s}_{1}(E)\wedge[x_{2}=0]=\tau_{*}(s_{1}(\tau^{*}E))=\tau_{*}[t=0]=[x_{1}=x_{2}=0]$. From [1, Example 11.2] we have that $s_{2}(E)=[x_{1}=x_{2}=0]$. Thus
@@
which we notice is different from $c_{2}(E)$. ∎

### v2

@@ 8. Remarks and examples
depending on how $k=k_{1}+\dots+k_{t}$ compares to $\kappa$}. Such a product is furthermore robust in the sense that it is commutative in the factors provided $k\leq{\kappa}$, which follows from Theorem 6.7 (vii), cf. Remark 6.8 (ii).
By Theorem 6.5 (iii), ${\bf s}(E,h+{\epsilon}h_{0})^{k}\mu=s(E\oplus\cdots\oplus E,h\oplus\cdots\oplus h+{\epsilon}(h_{0}\oplus\cdots\oplus h_{0})\wedge\mu$ for each $k$. Therefore $\lim_{{\epsilon}\to 0}c(E,h+{\epsilon}h_{0})\wedge\mu$ exists and is in ${\mathcal{QZ}}(X)$. Moreover it coincides with $c(E,h)\wedge\mu$ outside the degeneracy locus. However, in general it is not equal to ${\bf c}(E,h,h_{0})\wedge\mu$. This will be discussed in a forthcoming paper.
@@ 8.1. Difference between $c_{k}(E)$ and $c_{k}(E,Z)$
In general $c(E,Z)$ and $c(E)$ are different. The simplest example is when $Z=X$ and $Z^{\prime}\neq Z$. Then in general $s(E)$ will have a residue term, whereas $s(E,Z)=\omega$ and hence $c(E,Z)=c(E_{0})$.
@@
For instance, if $\tau\colon{\mathbb{C}}\to{\mathbb{C}}^{2}$, $t\mapsto(t,0)$, then ${\bf s}_{1}(\tau^{*}E)=[t=0]$. Hence, by Theorem 6.7 (iii), ${\bf s}_{1}(E)\wedge[x_{2}=0]=\tau_{*}(s_{1}(\tau^{*}E))=\tau_{*}[t=0]=[x_{1}=x_{2}=0]$. From [1, Example 11.2] we have that $s_{2}(E)=s_{2}(E,Z)=[x_{1}=x_{2}=0]$. Thus
@@
which is different from $c_{2}(E)$. ∎